SYSTEM OF LINEAR INEQUALITIES
SOLVE FOR X


4x+5>(>that sign has a line under it)2(x+1)

Answer by oscargut(2103)

Solution:
4x+5 > =2(x+1)
4x+5 > = 2x+2
4x-2x > = 2-5
2x > = -3
x > = -3/2
Answer: {x > = -3/2}